# Multiple-Frequencies Population-Based Training

### v1

@@ Abstract
Reinforcement Learning’s high sensitivity to hyperparameters is a source of instability and inefficiency, creating significant challenges for practitioners. Hyperparameter Optimization (HPO) algorithms have been developed to address this issue, among them Population-Based Training (PBT) stands out for its ability to generate hyperparameters schedules instead of fixed configurations. PBT trains a population of agents, each with its own hyperparameters, frequently ranking them and replacing the worst performers with mutations of the best agents. These intermediate selection steps can cause PBT to focus on short-term improvements, leading it to get stuck in local optima and eventually fall behind vanilla Random Search over longer timescales. This paper studies how this greediness issue is connected to the choice of evolution frequency, the rate at which the selection is done. We propose Multiple-Frequencies Population-Based Training (MF-PBT), a novel HPO algorithm that addresses greediness by employing sub-populations, each evolving at distinct frequencies. MF-PBT introduces a migration process to transfer information between sub-populations, with an asymmetric design to balance short and long-term optimization. Extensive experiments on the Brax suite demonstrate that MF-PBT improves sample efficiency and long-term performance, even without actually tuning hyperparameters.  [^1] Our code and experiments are available at: https://github.com/WaelDLZ/MF-PBT
@@ 2 Preliminaries
Hyperparameter Optimization (HPO, Bergstra et al. (2011); Feurer & Hutter (2019)) encompasses various approaches aimed at efficiently tuning hyperparameters to enhance performance and robustness of learning algorithms. Random Search Bergstra & Bengio (2012) is the baseline approach to HPO; the field then progressed towards more sophisticated techniques, notably meta-gradient methods Finn et al. (2017); Xu et al. (2018), sequential optimization Li et al. (2018); Falkner et al. (2018); Awad et al. (2021), and population-based approaches.
HPO methods generally follow one of two perspectives. Sequential optimization approaches treat HPO as the search for an optimal hyperparameter configuration within a fixed computational budget, assuming that once identified, this configuration can be reused across multiple training runs. In contrast, population-based methods integrate hyperparameter tuning into the training process itself, adapting hyperparameters dynamically to maximize the performance of a single model. In this work, we adopt the latter perspective, which justifies our exclusive focus on population-based methods in both our discussions and experiments.
@@ 2.1 Population-Based Training
Numerous extensions to PBT have been proposed, focusing on improving exploration and efficiency. Methods like PB2 Parker-Holder et al. (2020; 2021) use bandit theory to explore hyperparameter spaces, offering performance guarantees, particularly in small population settings. SEARL Franke et al. (2021) enhances sample efficiency in off-policy RL by employing a shared replay buffer across the population. BG-PBT Wan et al. (2022) integrates policy distillation Rusu et al. (2016) to jointly optimize neural architectures and hyperparameters.
@@ 5.2 Symmetric migration
We now assess the importance of the asymmetry in the migration process, which adds a protection against hyperparameter collapse by preventing greedy agents from corrupting steadier sub-populations. To test this, we compare MF-PBT with an alternative version where hyperparameters are always transferred along with weights, regardless of the $\delta$-values.
The training trajectories in Figure 5 show that while the asymmetry has little impact on Hopper, it yields improvements in most environments, particularly in the challenging Humanoid task. This indicates that the asymmetric design indeed enhances long-term performance.

### v2

@@ Abstract
Reinforcement Learning’s high sensitivity to hyperparameters is a source of instability and inefficiency, creating significant challenges for practitioners. Hyperparameter Optimization (HPO) algorithms have been developed to address this issue, among them Population-Based Training (PBT) stands out for its ability to generate hyperparameters schedules instead of fixed configurations. PBT trains a population of agents, each with its own hyperparameters, frequently ranking them and replacing the worst performers with mutations of the best agents. These intermediate selection steps can cause PBT to focus on short-term improvements, leading it to get stuck in local optima and eventually fall behind vanilla Random Search over longer timescales. This paper studies how this greediness issue is connected to the choice of evolution frequency, the rate at which the selection is done. We propose Multiple-Frequencies Population-Based Training (MF-PBT), a novel HPO algorithm that addresses greediness by employing sub-populations, each evolving at distinct frequencies. MF-PBT introduces a migration process to transfer information between sub-populations, with an asymmetric design to balance short and long-term optimization. Extensive experiments on the Brax suite demonstrate that MF-PBT improves sample efficiency and long-term performance, even without actually tuning hyperparameters.  [^1] Our code is available at: https://github.com/WaelDLZ/MF-PBT
@@ 2 Preliminaries
Hyperparameter Optimization (HPO, Bergstra et al. (2011); Feurer & Hutter (2019)) encompasses various approaches aimed at efficiently tuning hyperparameters to enhance performance and robustness of learning algorithms. Random Search Bergstra & Bengio (2012) is the baseline approach to HPO; the field then progressed towards more sophisticated techniques, notably meta-gradient methods Xu et al. (2018), sequential optimization Li et al. (2018); Falkner et al. (2018); Awad et al. (2021), and population-based approaches.
HPO methods generally follow one of two perspectives. Sequential optimization approaches treat HPO as the search for an optimal hyperparameter configuration within a fixed computational budget, assuming that once identified, this configuration can be reused across multiple training runs. In contrast, population-based methods integrate hyperparameter tuning into the training process, adapting hyperparameters dynamically to maximize the performance of a single model. In this work, we adopt the latter perspective, which justifies our exclusive focus on population-based methods in both our discussions and experiments.
@@ 2.1 Population-Based Training
Numerous extensions to PBT have been proposed, focusing on improving exploration and efficiency. Methods like PB2 Parker-Holder et al. (2020; 2021) use bandit theory to explore hyperparameter spaces, offering performance guarantees, particularly in small population settings. SEARL Franke et al. (2021) enhances sample efficiency in off-policy RL by employing a shared replay buffer across the population. AdaQN Vincent et al. (2025) further imrpoves sample efficiency in this setting by using the temporal-difference error as a fitness function. BG-PBT Wan et al. (2022) integrates policy distillation Rusu et al. (2016) to jointly optimize neural architectures and hyperparameters.
@@ 5.2 Symmetric migration
We now assess the importance of the asymmetric migration process, which enables information sharing between the sub-populations while adding a protection against hyperparameter collapse by preventing greedy agents from corrupting steadier sub-populations.
We compare MF-PBT to an alternative version where hyperparameters are always transferred along with weights, regardless of the $\delta$-values (Symmetric). To highlight the importance of migration, we also compare to a version that uses multiple-frequencies but no migration, consisting in 4 runs of 8-agents PBT.
The training trajectories in Figure 5 show the superiority asymmetric migration over the two alternatives, leading to better anytime and final performance.